How many copper (Cu) atoms are in 2.04 moles of copper?


What is the mass of 0.500 mol of manganese (Mn)?
Answer #1

1 mole Cu contains Avogadro's number 6.023 \times 10^{23} of atoms.

2.04 moles Cu will contain 2.04 \times 6.023 \times 10^{23} = 1.228 \times 10^{24} atoms.

The molar mass of Mn is 54.4 g/mol.

Hence, 1 mole of Mn will weigh 54.4 g

Hence, 0.500 moles of Mn will weigh \text { 0.500 mol} \times \text { 54.4 g/mol} = \text { 27.45 g}